# Think-on-Graph 3.0: Efficient and Adaptive LLM Reasoning on Heterogeneous Graphs via Multi-Agent Dual-Evolving Context Retrieval

### v1

@@ Abstract
Retrieval-Augmented Generation (RAG) and Graph-based RAG has become the important paradigm for enhancing Large Language Models (LLMs) with external knowledge. However, existing approaches face a fundamental trade-off. While graph-based methods are inherently dependent on high-quality graph structures, they face significant practical constraints: manually constructed knowledge graphs are prohibitively expensive to scale, while automatically extracted graphs from corpora are limited by the performance of the underlying LLM extractors, especially when using smaller, local-deployed models. This paper presents Think-on-Graph 3.0 (ToG-3), a novel framework that introduces Multi-Agent Context Evolution and Retrieval (MACER) mechanism to overcome these limitations. Our core innovation is the dynamic construction and refinement of a Chunk-Triplets-Community heterogeneous graph index, which pioneeringly incorporates a dual-evolution mechanism of Evolving Query and Evolving Sub-Graph for precise evidence retrieval. This approach addresses a critical limitation of prior Graph-based RAG methods, which typically construct a static graph index in a single pass without adapting to the actual query. A multi-agent system, comprising Constructor, Retriever, Reflector, and Responser agents, collaboratively engages in an iterative process of evidence retrieval, answer generation, sufficiency reflection, and, crucially, evolving query and subgraph. This dual-evolving multi-agent system allows ToG-3 to adaptively build a targeted graph index during reasoning, mitigating the inherent drawbacks of static, one-time graph construction and enabling deep, precise reasoning even with lightweight LLMs. Extensive experiments demonstrate that ToG-3 outperforms compared baselines on both deep and broad reasoning benchmarks, and ablation studies confirm the efficacy of the components of MACER framework.
footnotetext: Our code implementation are available at https://github.com/DataArcTech/RAG-Factory.
@@ 1 Introduction
Figure 2: Evolution of Retrieval-Augmented Generation Paradigms. (a) Naive RAG embeds raw documents and performs single-shot retrieval. (b) Graph-based RAG pre-builds a static graph once and retrieves from it. (c) ToG-3 introduces a four-agent loop—Retriever, Constructor, Reflector, Response—where the graph and the query sub-tasks co-evolve at runtime, yielding dynamic, query-adaptive context that converges to a minimal, sufficient subgraph.
Extensive experiments on complex multi-hop reasoning benchmarks demonstrate that our method achieves the highest average Exact Match and F1 scores on HotpotQA, 2WikiMultihopQA, and Musique. For broad reasoning tasks, ToG-3 also achieves remarkable win rates over baselines across comprehensiveness, diversity, and empowerment dimensions.
@@ 2.1 Graph-Based Retrieval-Augmented Generation
Recent advances in retrieval-augmented generation (RAG) have increasingly emphasized structural awareness to improve reasoning depth and contextual coherence. Edge et al. (2024) propose GraphRAG, which builds a knowledge graph (KG) from documents via LLM-based entity and relation extraction, then applies community detection to generate hierarchical summaries for global sensemaking. Guo et al. (2024) introduce LightRAG, which employs a dual-level retrieval system combining low-level fact retrieval and high-level semantic discovery using a compact KG, improving both efficiency and coverage. Further building on this idea, Gutiérrez et al. (2024; 2025) present a non-parametric continual learning framework that uses Personalized PageRank over an open KG to enable associative, multi-hop reasoning. Other structure-augmented RAG methods include RAPTOR (Sarthi et al., 2024), Chen et al. (2023) enhance sense-making but often introduce noise through uncontrolled summarization or lack explicit support for multi-hop reasoning.
@@ 2.2 Knowledge Graphs in RAG and Hybrid Approaches
The integration of structured knowledge into LLM reasoning has long been pursued to improve faithfulness and interpretability. Early KG-augmented RAG systems retrieve triples from static external knowledge bases such as Wikidata or Freebase to ground model outputs (Sun et al., 2023). However, these sources are often incomplete, outdated, or misaligned with domain-specific content. To overcome this, hybrid RAG frameworks (Ma et al., 2024) combine unstructured text and structured KGs to balance breadth and precision. Chain-of-Knowledge (CoK) (Li et al., 2024) retrieves from multiple structured sources including Wikipedia, Wikidata, and Wikitable to ground LLM responses. HybridRAG (Sarmah et al., 2024) fuses vector-based and KG-based retrievers, demonstrating superior reasoning performance compared to either modality alone.
@@ 2.3 Iterative and Reflective Reasoning in LLMs
Enabling LLMs to reason iteratively has been shown to improve accuracy and faithfulness. ITER-RETGEN (Shao et al., 2023) introduces an iterative loop that alternates between retrieval and generation, using generated hypotheses to guide further search. Trivedi et al. (2023) combine Chain-of-Thought (CoT) with retrieval, interleaving reasoning steps with evidence gathering, significantly improving performance on multi-hop QA. Self-RAG (Asai et al., 2023) equips LLMs with reflection tokens to decide when to retrieve and whether the output is hallucinated. ReAct (Yao et al., 2023a) combines reasoning traces with external actions, enabling task decomposition and environment interaction. Other efforts focus on continual learning for LLMs, where RAG serves as a non-parametric alternative to fine-tuning (Shi et al., 2024). Continual pretraining (Jin et al., 2022) and instruction tuning (Zhang et al., 2023) can update model parameters but suffer from catastrophic forgetting (Huang et al., 2024). Model editing methods (Yao et al., 2023b) offer fine-grained updates but struggle with generalization.
@@ 3.2.2 Offline Index Construction
Algorithm 1 in Appendix. B details the one-time construction of the universal index $\mathcal{G}$. A key design choice is the use of a single frozen encoder $E_{\theta}$ (e.g., jina-mebedding-v3 (Sturua et al., 2024)) to embed all nodes—regardless of type—into a unified 1024-dimensional dense vector space. This enables efficient vector search across all node types during retrieval.
@@ 3.3 The MACER Process: Multi-Agent Context Evolution and Retrieval
Figure 3: Multi-Agent Dual-Evolving Context Retrieval-Response Loop. The Retriever fetches an initial chunk–triplet–community subgraph. The Response Agent produces an answer; the Reflector Agent judges sufficiency (reward=1/0). If insufficient (reward=0), the Reflector evolves the query into sub-queries while the Constructor evolves the subgraph (sub-graph refinement). The loop repeats until the context becomes sufficient or the horizon is reached, after which the Response Agent synthesizes the final answer from the full trajectory.
@@
: At each step $k$, the state $s_{k}=(q,\mathcal{G}_{k},\mathcal{H}_{k})$ captures the complete reasoning context, including the original query $q$, the current evidence subgraph $\mathcal{G}_{k}$ retrieved by Retriever Agent $\pi_{\text{ret}}$, and the trajectory history $\mathcal{H}_{k}={(q^{\prime}_{i},a_{i},r_{i},\mathcal{G}i)}_{i=0}^{k-1}$ of all previous sub-queries, answers, rewards, and sub-graphs.
@@
$r_{k}=\begin{cases}1&\text{if }\texttt{Suff}(q,\mathcal{G}_{k},a_{k})=1\quad\text{(sufficient context)}\\ 0&\text{otherwise.}\end{cases}$ (2)
This reward signal directly dictates the termination condition and guides the refinement process.
@@
The complete MACER process, now cast as an MDP, is summarized in Algorithm 2. The loop continues until the Reflector’s policy $\pi_{\text{ref}}$ outputs the stop action (via $r_{k}=1$) or a maximum horizon $K$ is reached. The final answer $a^{*}$ is synthesized from the full trajectory $\mathcal{H}_{k}$ of states and actions, ensuring faithfulness to the evolved evidence. This MDP formulation provides the formal foundation for establishing the convergence of the MACER process under mild assumptions, as detailed in Appendix. I. This iterative refinement allows ToG-3 to start from a potentially weak initial graph but specialize it towards the reasoning path of the specific query, converging on a high-quality evidence subgraph $\mathcal{G}^{*}_{q}$. This evolving and refinement mechanism alleviate the three fundamental weaknesses of small LMs in static GraphRAG, including incomplete triplet recall, insufficient knowledge details and high parsing failure of LLMs’ output, as mentioned in Section 1.
@@ 4.1 Experimental Setup
To comprehensively evaluate the reasoning capabilities of RAG systems, we conduct experiments on two distinct categories of tasks: Deep Reasoning Tasks including HotpotQA (Yang et al., 2018), 2WikiMultiHopQA (Ho et al., 2020) and Musique (Trivedi et al., 2022) and Broad Reasoning Tasks including 4 subsets of UltraDomain (Qian et al., 2025) benchmark. Detailed statistics for all datasets are provided in Table 4 and Appendix. C.
@@
For Deep Reasoning Tasks, we follow standard QA evaluation practices with Exact Match (EM) and F1 Score. For Broad Reasoning Tasks, we adopt a multi-dimensional LLM-based evaluation approach including Comprehensiveness, Diversity and Empowerment following (Guo et al., 2024). Metrics detail are provide Appendix.E.
@@
We compare ToG-3 against the following state-of-the-art RAG methods across all datasets, including NaiveRAG (Gao et al., 2023), ToG-2 (Ma et al., 2024), GraphRAG (Edge et al., 2024), LightRAG (Guo et al., 2024), MiniRAG (Fan et al., 2025) and HippoRAG-2 (Gutiérrez et al., 2025). Baselines details can be found in Appendix.D. For graph-based methods, we maintain identical chunk sizes (1024 tokens) and use the same LLM (Qwen2.5-32B-Instruct (Yang et al., 2024)) for all extraction and generation tasks to eliminate model capability variations. Implementation details are provide Appendix.A.
@@ 4.2 Result of Deep Reasonging Benchmark
Results shown in Table 1 represent the average of three independent reasoning experiments. Previsou Graph-based methods like GraphRAG that rely on LLM-based graph construction show limited performance. Their performance is the lowest, particularly in terms of F1 scores as shown in Figure 4(b), which can be attributed to a lack of focus on deep factual reasoning and a tendency to produce verbose responses, resulting in low token-level recall. More detailed precision and recall results are provided in Appendix. F.1. ToG-2, without leveraging well-curated knowledge graphs like Freebase and Wikidata, demonstrates moderate performance in open-domain settings. NaiveRAG achieves competitive third-place results by avoiding graph construction limitations and relying solely on retrieved documents for response generation. HippoRAG-2 emerges as the strongest baseline, employing an efficient embedding model with Personalized PageRank algorithm and LLM-based triple filtering to achieve second-best performance. However, our proposed method consistently outperforms all competitors, achieving the highest average EM (0.453) and F1 (0.312) scores across all three benchmarks. This superior performance is attributed to our novel Chunk-Triplets-Community heterogeneous graph architecture and the Multi-Agent Context Evolution and Retrieval (MACER) framework, which enables adaptive subgraph refinement and evolving query decomposition for complex reasoning tasks and overcomes the graph construction challenges that plague other graph-based RAG systems.
@@
As shown in Figure 4, the average performance of the baselines and our method across the HotpotQA, 2WikiMultiHopQA, and Musique datasets generally follows a descending trend. This pattern can be attributed to the following reasons: HotpotQA (Yang et al., 2018): Although widely used, this dataset has been shown to provide a weaker test of multi-hop reasoning due to the presence of numerous spurious cues and shortcut signals (Trivedi et al., 2022; Gutiérrez et al., 2024). Musique (Trivedi et al., 2022): A challenging multi-hop QA dataset comprising approximately requiring 2–4 hops, which emphasizes a comprehensive evaluation of multi-step reasoning abilities. Musique is designed to feature diverse and complex reasoning paths, necessitating the integration of information across multiple hops to arrive at correct answers.
@@ 4.3 Result of Broad Reasoning Tasks
As shwon in Figure 5, The four heatmaps clearly demonstrate that the five methods can be distinctly divided into two clusters: the upper-right region (predominantly red, indicating superior performance) and the lower-left region (predominantly blue, indicating inferior performance). Specifically, ToG-3, GraphRAG, and LightRAG exhibit significantly higher win rates compared to NaiveRAG and HippoRAG-2. Detailed win rates (%) of baselines v.s. ToG-3 across four datasets are provided in Table 6 of Appendix. F. Our framework outperforms NaiveRAG by substantial margins (up to 88.8% overall win rate on Legal dataset and 72.9% average win rate on all four datasets), highlighting the limitations of chunk-based retrieval for complex queries. While GraphRAG shows competitive performance in comprehensiveness due to its extensive community summarization and retrival, ToG-3 achieves better balance across all metrics, particularly excelling in diversity and empowerment through its heterogeneous graph architecture that integrates chunk-level, triplet-level, and community-level information. Detailed ELO rating calculation for broad reasoning tasks can be found in Appendix. F.3. The multi-agent dual-evolving context retrieval mechanism enables both deep knowledge reasoning through entity-relation exploration and broad community reasoning. This balanced architectural approach makes ToG-3 particularly effective for real-world applications requiring both comprehensive coverage and precise, actionable insights. Our analysis reveals that, on average, 20% of the samples require one evolving-context iteration, 32% require two iterations, and 48% require three iterations. Case studies of ToG-3 retrieval and response output are provided in Appendix. G.
4.4 Analysis of Computation Cost
Dataset Method Graph Statistics Indexing Inference Avg.
Entities Relations Communities Time (h) Time (s/q) EM
HotpotQA ToG-3 37,358 30,987 5,041 12.5 16.82 0.639
GraphRAG 94,376 73,265 10,981 15.8 8.91 0.337
LightRAG 94,578 76,157 - 12.1 6.54 0.308
2WikiMultihopQA ToG-3 19,311 21,077 3,417 8.2 14.95 0.500
GraphRAG 50,556 37,840 6,261 10.3 7.45 0.439
LightRAG 50,177 37,995 - 7.8 5.23 0.420
Musique ToG-3 32,842 39,134 6,258 9.7 13.24 0.221
GraphRAG 106,042 83,139 9,407 13.2 9.37 0.109
LightRAG 94,621 75,923 - 10.3 7.12 0.082
Average ToG-3 29,837 30,399 4,905 10.13 15.00 0.453
GraphRAG 83,658 64,748 8,883 13.10 8.58 0.295
LightRAG 79,792 63,358 - 10.06 6.30 0.270
The Table 2 reveal a consistent accuracy-efficiency trade-off across all datasets. We observed that during the indexing phase, GraphRAG required the longest processing time, averaging 13.10 hours. This is primarily due to its need to extract a large number of triplets and generate community summaries. In comparison, both ToG-3 and LightRAG showed similar indexing times, at 10.13 and 10.06 hours respectively. Although ToG-3 also involves community summary generation, it constructs the graph more efficiently by extracting fewer relational structures during graph initialization compared to both LightRAG and GraphRAG. While LightRAG achieve faster inference times, they suffer from lower accuracy due to redundant graph elements or simpler retrieval mechanisms. GraphRAG’s expensive two-stage indexing yields suboptimal results despite longer processing times. ToG-3 demonstrates an effective balance: its efficient heterogeneous graph construction produces refined knowledge bases across all datasets, and while its multi-agent reasoning requires higher inference time, this cost is directly justified by its best performance on all benchmarks, making it ideal for quality-sensitive applications requiring reliable reasoning capabilities. Note that HippoRAG-2 was excluded from the comparative analysis due to its reliance on OpenIE-based extraction rather than pure LLM extraction, which differs fundamentally from the approaches under investigation.

### v2

@@ Abstract
Graph-based Retrieval-Augmented Generation (GraphRAG) has become the important paradigm for enhancing Large Language Models (LLMs) with external knowledge. However, existing approaches are constrained by their reliance on high-quality knowledge graphs: manually built ones are not scalable, while automatically extracted ones are limited by the performance of LLM extractors, especially when using smaller, local-deployed models. To address this, we introduce Think-on-Graph 3.0 (ToG-3), a novel framework featuring a Multi-Agent Context Evolution and Retrieval (MACER) mechanism. Its core contribution is the dynamic construction and iterative refinement of a Chunk-Triplets-Community heterogeneous graph index, powered by a Dual-Evolution process that adaptively evolves both the query and the retrieved sub-graph during reasoning. ToG-3 dynamically builds a targeted graph index tailored to the query, enabling precise evidence retrieval and reasoning even with lightweight LLMs. Extensive experiments demonstrate that ToG-3 outperforms compared baselines on both deep and broad reasoning benchmarks, and ablation studies confirm the efficacy of the components of MACER framework. The source code are available in https://github.com/DataArcTech/ToG-3.
11footnotetext: Equal Contribution22footnotetext: Corresponding Author
@@ 1 Introduction
Figure 2: Evolution of Retrieval-Augmented Generation Paradigms. (a) Naive RAG embeds raw documents and performs single-shot retrieval. (b) Graph-based RAG pre-builds a static graph once and retrieves from it. (c) ToG-3 introduces a four-agent loop—Retriever, Constructor, Reflector, Reranker, Responser—where the graph and the query sub-tasks co-evolve at runtime, yielding dynamic, query-adaptive context that converges to a minimal, sufficient subgraph.
@@ 2.1 Graph-Based Retrieval-Augmented Generation
Recent advances in retrieval-augmented generation (RAG) have increasingly emphasized structural awareness to improve reasoning depth and contextual coherence. edge2024local propose GraphRAG, which builds a knowledge graph (KG) from documents via LLM-based entity and relation extraction, then applies community detection to generate hierarchical summaries for global sensemaking. guo2024lightrag introduce LightRAG, which employs a dual-level retrieval system combining low-level fact retrieval and high-level semantic discovery using a compact KG, improving both efficiency and coverage. Further building on this idea, hipporag; gutiérrez2025HippoRAG2 present a non-parametric continual learning framework that uses Personalized PageRank over an open KG to enable associative, multi-hop reasoning. Other structure-augmented RAG methods include RAPTOR (raptor), chen2023walkingmemorymazecontext enhance sense-making but often introduce noise through uncontrolled summarization or lack explicit support for multi-hop reasoning. Note that RL-based frameworks such as GraphRAG-R1 (yu2025graphrag-r1) and Graph-R1 (luo2025graph-r1) utilize existing Graph-based RAG methods as their retrieval components and train an end-to-end agentic framework. Our work, in contrast, proposes a complementary approach to improve the underlying retrieval paradigm itself. Consequently, ToG-3 could also serve as a plug-in component to enhance such RL frameworks.
@@ 2.2 Knowledge Graphs in RAG and Hybrid Approaches
The integration of structured knowledge into LLM reasoning has long been pursued to improve faithfulness and interpretability. Early KG-augmented RAG systems retrieve triples from static external knowledge bases such as Wikidata or Freebase to ground model outputs (sun2023tog). However, these sources are often incomplete, outdated, or misaligned with domain-specific content. To overcome this, hybrid RAG frameworks (ma2024tog2) combine unstructured text and structured KGs to balance breadth and precision. Chain-of-Knowledge (CoK) (li2024cok) retrieves from multiple structured sources including Wikipedia, Wikidata, and Wikitable to ground LLM responses. HybridRAG (sarmah2024hybridrag) fuses vector-based and KG-based retrievers, demonstrating superior reasoning performance compared to either modality alone.
@@ 2.3 Iterative and Reflective Reasoning in LLMs
Enabling LLMs to reason iteratively has been shown to improve accuracy and faithfulness. ITER-RETGEN (shao2023enhancingretrievalaugmentedlargelanguage) introduces an iterative loop that alternates between retrieval and generation, using generated hypotheses to guide further search. trivedi2023interleavingretrievalchainofthoughtreasoning combine Chain-of-Thought (CoT) with retrieval, interleaving reasoning steps with evidence gathering, significantly improving performance on multi-hop QA. Self-RAG (asai2023selfraglearningretrievegenerate) equips LLMs with reflection tokens to decide when to retrieve and whether the output is hallucinated. ReAct (yao2023react) combines reasoning traces with external actions, enabling task decomposition and environment interaction. Other efforts focus on continual learning for LLMs, where RAG serves as a non-parametric alternative to fine-tuning (shi2024continual). Continual pretraining (lifelong) and instruction tuning (citb) can update model parameters but suffer from catastrophic forgetting (huang24mitigating). Model editing methods (yao23editing) offer fine-grained updates but struggle with generalization.
@@ 3.2.2 Offline Index Construction
Algorithm 1 in Appendix. B details the one-time construction of the universal index $\mathcal{G}$. A key design choice is the use of a single frozen encoder $E_{\theta}$ (e.g., jina-mebedding-v3 (sturua2024jinaembeddingsv3)) to embed all nodes—regardless of type—into a unified 1024-dimensional dense vector space. This enables efficient vector search across all node types during retrieval.
@@ 3.3 The MACER Process: Multi-Agent Context Evolution and Retrieval
Figure 3: Multi-Agent Dual-Evolving Context Retrieval-Response Loop. The Retriever fetches an initial chunk–triplet–community subgraph and the Reranker reranks and selects the top-n most relevant pieces of evidence.. The Response Agent produces an answer; the Reflector Agent judges sufficiency (reward=1/0). If insufficient (reward=0), the Reflector evolves the query into sub-queries while the Constructor evolves the subgraph (sub-graph refinement). The loop repeats until the context becomes sufficient or the horizon is reached, after which the Response Agent synthesizes the final answer from the full trajectory.
@@
: At each step $k$, the state $s_{k}=(q,\mathcal{G}_{k},\mathcal{H}_{k})$ captures the complete reasoning context, including the original query $q$, the current evidence subgraph $\mathcal{G}_{k}$ retrieved by Retriever Agent $\pi_{\text{ret}}$ and reranked by Reranker Agent $\pi_{\text{rer}}$, and the trajectory history $\mathcal{H}_{k}={(q^{\prime}_{i},a_{i},r_{i},\mathcal{G}i)}_{i=0}^{k-1}$ of all previous sub-queries, answers, rewards, and sub-graphs.
@@
$r_{k}=\begin{cases}1&\text{if }\texttt{Suff}(q,\mathcal{G}_{k},a_{k})=1\\ 0&\text{otherwise.}\end{cases}$ (2)
This reward signal is produced by the Reflector Agent to determine if the current context evidence is sufficient to answer the user’s query.
@@
The complete MACER process, now cast as an MDP, is summarized in Algorithm 2. The loop continues until the Reflector’s policy $\pi_{\text{ref}}$ outputs the stop action (via $r_{k}=1$) or a maximum horizon $K$ is reached. The final answer $a^{*}$ is synthesized from the full trajectory $\mathcal{H}_{k}$ of states and actions, ensuring faithfulness to the evolved evidence. This MDP formulation provides the formal foundation for establishing the convergence of the MACER process under mild assumptions, as detailed in Appendix. K. This iterative refinement allows ToG-3 to start from a potentially weak initial graph but specialize it towards the reasoning path of the specific query, converging on a high-quality evidence subgraph $\mathcal{G}^{*}_{q}$. This evolving and refinement mechanism alleviate the three fundamental weaknesses of small LMs in static GraphRAG, including incomplete triplet recall, insufficient knowledge details and high parsing failure of LLMs’ output, as mentioned in Section 1.
@@ 4.1 Experimental Setup
To comprehensively evaluate the reasoning capabilities of RAG systems, we conduct experiments on two distinct categories of tasks: Deep Reasoning Tasks including HotpotQA (yang2018hotpotqa), 2WikiMultiHopQA (2wiki) and Musique (musique) and Broad Reasoning Tasks including 4 subsets of UltraDomain (qian2025memorag) benchmark. Detailed statistics for all datasets are provided in Table 3 and Appendix. C.
@@
For Deep Reasoning Tasks, we follow standard QA evaluation practices with Exact Match (EM)(Following ToG and ToG-2 (sun2023tog; ma2024tog2), we employ a substring-based Exact Match metric.) and F1 Score. For Broad Reasoning Tasks, we adopt a multi-dimensional LLM-based evaluation approach including Comprehensiveness, Diversity and Empowerment following (guo2024lightrag). Metrics detail are provide Appendix.E.
@@
We compare ToG-3 against the following state-of-the-art RAG methods across all datasets, including NaiveRAG (gao2023rag_survey), ToG-2 (ma2024tog2), GraphRAG (edge2024local), LightRAG (guo2024lightrag), MiniRAG (fan2025minirag) and HippoRAG-2 (gutiérrez2025HippoRAG2). Baselines details can be found in Appendix.D. For graph-based methods, we maintain identical chunk sizes (1024 tokens) and use the same LLM (Qwen2.5-32B-Instruct (qwen2.5)) for all extraction and generation tasks to eliminate model capability variations. Implementation details are provide Appendix.A.
@@ 4.2 Result of Deep Reasonging Benchmark
Results shown in Table 1 represent the average of three independent reasoning experiments. Previsou Graph-based methods like GraphRAG that rely on LLM-based graph construction show limited performance. Their performance is the lowest, particularly in terms of F1 scores as shown in Figure 4(b), which can be attributed to a lack of focus on deep factual reasoning and a tendency to produce verbose responses, resulting in low token-level recall. More detailed precision and recall results are provided in Appendix. F.1. ToG-2, without leveraging well-curated knowledge graphs like Freebase and Wikidata, demonstrates moderate performance in open-domain settings. NaiveRAG achieves competitive third-place results by avoiding graph construction limitations and relying solely on retrieved documents for response generation. HippoRAG-2 emerges as the strongest baseline, employing an efficient embedding model with Personalized PageRank algorithm and LLM-based triple filtering to achieve second-best performance. However, our proposed method consistently outperforms all competitors, achieving the highest average EM (0.474) and F1 (0.345) scores across all three benchmarks. This superior performance is attributed to our novel Chunk-Triplets-Community heterogeneous graph architecture and the Multi-Agent Context Evolution and Retrieval (MACER) framework, which enables adaptive subgraph refinement and evolving query decomposition for complex reasoning tasks and overcomes the graph construction challenges that plague other graph-based RAG systems. Additional Baselines are provided in Appendix.H.
@@
As shown in Figure 4, the average performance of the baselines and our method across the HotpotQA, 2WikiMultiHopQA, and Musique datasets generally follows a descending trend. This pattern can be attributed to the following reasons: HotpotQA (yang2018hotpotqa): Although widely used, this dataset has been shown to provide a weaker test of multi-hop reasoning due to the presence of numerous spurious cues and shortcut signals (musique; hipporag). Musique (musique): A challenging multi-hop QA dataset comprising approximately requiring 2–4 hops, which emphasizes a comprehensive evaluation of multi-step reasoning abilities. Musique is designed to feature diverse and complex reasoning paths, necessitating the integration of information across multiple hops to arrive at correct answers.
@@ 4.3 Result of Broad Reasoning Tasks
As shwon in Figure 5, The four heatmaps clearly demonstrate that the five methods can be distinctly divided into two clusters: the upper-right region (predominantly red, indicating superior performance) and the lower-left region (predominantly blue, indicating inferior performance). Specifically, ToG-3, GraphRAG, and LightRAG exhibit significantly higher win rates compared to NaiveRAG and HippoRAG-2. Detailed win rates (%) of baselines v.s. ToG-3 across four datasets are provided in Table 5 of Appendix. F. Our framework outperforms NaiveRAG by substantial margins (up 75.0% average win rate on all four datasets), highlighting the limitations of chunk-based retrieval for complex queries. While GraphRAG shows competitive performance in comprehensiveness due to its extensive community summarization and retrival, ToG-3 achieves better balance across all metrics, particularly excelling in diversity and empowerment through its heterogeneous graph architecture that integrates chunk-level, triplet-level, and community-level information. Detailed ELO rating calculation for broad reasoning tasks can be found in Appendix. F.3. The multi-agent dual-evolving context retrieval mechanism enables both deep knowledge reasoning through entity-relation exploration and broad community reasoning. Our analysis reveals that, on average, 20% of the samples require one evolving-context iteration, 32% require two iterations, and 48% require three iterations.
Table 2: Ablation studies of MACER components and foundation model scaling. Standard ToG-3 settings incorporates all MACER components, employs the Qwen2.5-32B-instruct as the backbone LLM, and utilizes the Jina-v3-embedding model for representation encoding and Jina-reranker-v2 for reranking the retrieved evidence.
Ablation Setting HotpotQA 2WikiMultihopQA Musique Average
EM F1 EM F1 EM F1 EM F1
MACER Components Ablation
w/o Evolving Query 0.614 0.495 0.440 0.227 0.198 0.141 0.417 0.288
w/o Evolving Sub-Graph 0.629 0.525 0.486 0.258 0.223 0.158 0.446 0.314
w/o Community Node 0.656 0.572 0.514 0.283 0.236 0.169 0.469 0.341
Foundation Model Scaling Abalation
LLM Model
Qwen2.5-14B 0.587 0.521 0.480 0.255 0.218 0.154 0.428 0.310
Qwen2.5-72B 0.683 0.592 0.550 0.305 0.255 0.182 0.496 0.360
Embedding Model
Qwen3-Embed-0.6B 0.653 0.571 0.532 0.294 0.244 0.176 0.476 0.347
Qwen3-Embed-4B 0.658 0.577 0.535 0.296 0.247 0.179 0.480 0.351
Detailed comparison of time and token Consumption across different methods are provided in Appendix.G. Case studies of ToG-3 retrieval and response output are provided in Appendix. I.